# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v5

@@ 1 Introduction
Differential Privacy (DP) has become a foundational framework for safeguarding individual privacy while enabling meaningful data analysis [Dwo06]. In real-world applications, Local Differential Privacy (LDP) is widely adopted as it eliminates the need for a trusted curator by applying noise to each user’s data before aggregation [CJK+18, DKY17, YGZ+24, DPT17]. However, this decentralized approach often results in significant utility loss due to excessive noise.
@@
To address this trade-off, the shuffle model introduces a trusted shuffler between users and the aggregator [CSU+19, EFM+19, BEM+17]. This work considers the single-message shuffle model, where each client sends a report of their data and these reports are then anonymized and randomly shuffled before being sent to the server [FMT23a]. Shuffle DP improves the privacy-utility trade-off while minimizing trust assumptions to a single shuffler, making it a promising model for real-world deployment [WDX+20, IMC22, CZ21]. For instance, privacy amplification by shuffling was used in Apple and Google’s Exposure Notification Privacy-preserving Analytics [AG21].
@@
The “amplification-by-shuffling” theorem in the shuffle model implies that when each of the $n$ users randomizes their data using an $\varepsilon_{0}$-LDP mechanism, the collection of shuffled reports satisfies $(\varepsilon(\varepsilon_{0},\delta,n),\delta)$-DP, where $\varepsilon(\varepsilon_{0},\delta,n)\ll\varepsilon_{0}$ for sufficiently large $n$ and not too small $\delta$ [FMT23a]. A central theoretical challenge is to characterize and compute this privacy amplification effect. A tighter bound allows for a larger value of $\varepsilon_{0}$ while still achieving $(\varepsilon,\delta)$-DP after shuffling, thereby improving the utility of the mechanism. It also enables a more accurate analysis of the overall $(\varepsilon,\delta)$-guarantees resulting from shuffling $n$ independent $\varepsilon_{0}$-LDP reports.
@@
Researchers are interested in two types of upper bounds on privacy amplification. The first is a generic bound, which provides a uniform guarantee for all $\varepsilon_{0}$-DP local randomizers. The second is a specific bound, which is tailored for a given $\varepsilon_{0}$-DP local randomizer. Generic bounds are primarily studied from a theoretical perspective, as they allow the derivation of asymptotic expressions for $\varepsilon(\varepsilon_{0},n,\delta)$ [EFM+19, BBGN19, FMT22]. In contrast, specific bounds are usually much tighter and are of significant practical importance, as they provide precise guarantees for concrete mechanisms used in real-world applications.
@@
Among the various strategies proposed to analyze this effect [CSU+19, EFM+19, BKM+20], decomposition-based approaches are the most popular [BBGN19, FMT22, FMT23a]. Two prominent frameworks are the privacy blanket by Balle et al. [BBGN19] and the clone paradigm by Feldman et al., which includes the standard clone [FMT22] and the stronger clone [FMT23a]. These approaches rely on decomposing the probability distributions induced by a local randomizer under different inputs. A decomposition naturally leads to a reduction, which yields an upper bound on the privacy amplification via shuffling.
@@
The privacy blanket framework designs a tailored decomposition to obtain a bound for each specific local randomizer [BBGN19]. However, this bound cannot be computed precisely in its original form. To address this, the framework resorts to further approximations—such as Hoeffding’s and Bennett’s inequalities—which yield looser but computable bounds. A computable generic bound is also derived in a similar manner.
@@
The standard clone adopts a simple, unified decomposition that yields a generic bound applicable to all $\varepsilon_{0}$-DP local randomizers [FMT22]. This bound can be computed precisely via a dedicated numerical algorithm, and empirically outperforms the computable generic bound derived from the privacy blanket framework. However, the standard clone does not provide specific bounds, except in the special case of $k$-ary Randomized Response.
@@
The stronger clone was introduced with a more refined decomposition and was expected to yield improved bounds for both generic and specific local randomizers [FMT23a]. Unfortunately, a critical flaw was later identified in the proof’s core lemma. A corrected version was released on arXiv [FMT23b], which showed that the original generic bounds hold only for a restricted class of local randomizers. The specific bounds were also revised and replaced with a version that is much weaker and lacks an efficient computation method.
@@ 1.1 Our contributions
In prior work on Shuffle DP, the specific bounds of the blanket framework has seen limited use; instead, its associated analysis has mostly been applied in specific cases when the blanket distribution is uniform, as in $k$-ary randomized response ($k$-RR) [BBGN19], optimal local hash [WDX+20] or in multi-message shuffle models where uniform blanket noise is directly injected [LWY22]. Although the clone and blanket appear somewhat similar, no formal connection between them has been established.
@@
These two definitions are interconvertible: treating $(D_{1},D_{2},\ldots,D_{k})$ as a single data $D$ yields joint composition as a special case of $k$-fold composition; conversely, setting $D_{1}=D_{2}=\cdots=D_{k}=D$ makes $k$-fold composition a special case of joint composition. Our analysis of joint composition also applies to the $k$-fold composition setting. However, the joint composition framework captures a more general scenario—namely, when $D_{1},D_{2},\dots,D_{k}$ are not necessarily identical and may even belong to different domains. It is important to note that our setting concerns the composition of LDP mechanisms in the single-message shuffle model, in which $D$ or the tuple $(D_{1},D_{2},\dots,D_{k})$ corresponds to the data held by an individual user. Moreover, during the shuffling process, these outputs are treated as a single tuple and are shuffled together as a unit.
Joint composition is widely used in LDP applications such as joint distribution estimation and heavy hitter detection [DFSC22, RYY+18, BS15, Kik22]. For example, when each user’s data contains $d$ attributes, applying an $\frac{\varepsilon_{0}}{d}$-LDP mechanism to each attribute ensures overall $\varepsilon_{0}$-LDP while preserving inter-attribute correlations. Our experimental results show that existing methods yield relatively loose bounds in this setting, whereas our algorithm computes significantly tighter results by leveraging the optimal bounds.
@@
It is particularly suited for scenarios in which the analyst aims to compute multiple statistical queries—each corresponding to a distinct local mechanism $\mathcal{R}_{1},\dots,\mathcal{R}_{m}$. Parallel composition partitions the user population into $m$ non-overlapping subsets, assigning each subset to one of the $m$ estimation tasks. As a result, each subset can utilize the full privacy budget $\varepsilon_{0}$, rather than splitting the budget across all tasks. This technique has been widely adopted in the literature [BNST17, WXY+19, WDZ+19, WLJ21, CKS18] to achieve better utility compared to naive budget division.
It is worth noting that parallel composition also characterizes Poisson subsampling in the shuffle model [BBG18, Ste22, BBG20]: each user participates in a mechanism $\mathcal{R}^{*}$ with probability $p^{*}$, which is equivalent to the parallel composition $\mathcal{R}_{[1,2]}^{\parallel,\boldsymbol{p}}(x)$, where $\mathcal{R}_{1}(x)=\mathcal{R}^{*}(x)$, $\mathcal{R}_{2}(x)=\bot$, and $p_{1}=p^{*}$, $p_{2}=1-p^{*}$.
To compute the privacy amplification of parallel composition in the shuffle model was previously studied in [WPL+24], but their analysis inherited a technical flaw from [FMT23a], resulting in incorrect conclusions. Our method provides the first correct computation of the optimal bound in this setting.
@@ 1.2 Related Work
We now discuss prior work most relevant to the technical contributions of our paper. Koskela et al. [KHH23] investigate $k$-fold composition of Shuffle DP protocols. Their approach computes the Privacy Loss Random Variable (PLRV) of a dominating pair for a single-round shuffle protocol, and then applies the FFT to analyze multiple rounds (i.e., composition over $k$ executions) [KJH20]. In contrast, our work focuses on joint composition, where a single-round shuffle is applied to a composite local randomizer constructed from multiple sub-randomizers.
Some subtle aspects warrant further clarification, particularly regarding how coupling among sub-randomizers affects the composition analysis. In the single-message shuffle model, when the local randomizer is a joint composition, the outputs of the sub-local randomizers for a single user are inherently coupled. From a theoretical standpoint, privacy analysis in this setting ($\mathcal{S}\circ(\mathcal{M}_{1}\times\mathcal{M}_{2}\times\dots\times\mathcal{M}_{k})^{n}$) is not equivalent to applying composition to the individual mechanisms $\mathcal{R}_{i}$ after shuffling $(\mathcal{S}_{1}\circ\mathcal{M}_{1}^{n},\mathcal{S}_{2}\circ\mathcal{M}_{2}^{n},\dots,\mathcal{S}_{k}\circ\mathcal{M}_{k}^{n})$. Furthermore, the standard compositional properties of the PLRV do not apply.
In [KHH23], computing the PLRV requires a dominating pair [ZDW22], obtained either from the stronger clone (in the general case) or from the privacy blanket (in the case of $k$-RR). In the general case, [KHH23] inherits a flaw that originates from the original version of the stronger clone construction (see Section 3.3). For any specific local randomizer, our work demonstrates that the privacy bound derived from the blanket framework is strictly tighter than that obtained via the clone.
Theoretically, computing the PLRV of $(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$ (as defined in Section 3.2) and applying the Hockey-Stick divergence $D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}}\parallel P_{1}^{\mathsf{B}})$ yields the same amplification bound as computing it directly from the GPARV; both approaches evaluate $D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}}\parallel P_{1}^{\mathsf{B}})$. However, for general local randomizers, the PLRV of $(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$ is analytically intractable due to the complex structure of the blanket distribution. GPARV overcomes this limitation by expressing the privacy loss in a convolution-friendly form, enabling efficient FFT-based evaluation even in the absence of a closed-form expression for the PLRV.
@@ 1.3 Roadmap
Section 5 introduces the Generalized Privacy Amplification Random Variable (GPARV) and an FFT-based algorithm for computing the optimal privacy amplification bound. We also clarify the relationship between GPARV and the original PARV introduced by Balle et al. [BBGN19]. Section 6 presents our experimental results.
@@ 2 Preliminaries
Following conventions in the shuffle model based on randomize-then-shuffle [BBGN19, CSU+19], we define a single-message protocol $\mathcal{P}$ in the shuffle model as a pair of algorithms $\mathcal{P}=(\mathcal{R},\mathcal{A})$, where $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$, and $\mathcal{A}:\mathbb{Y}^{n}\to\mathbb{O}$. We call $\mathcal{R}$ the local randomizer, $\mathbb{Y}$ the message space of the protocol, $\mathcal{A}$ the analyzer, and $\mathbb{O}$ the output space.
@@ 3 Review of Existing Analysis Techniques
In this section, we review existing analysis techniques for studying privacy amplification in the shuffle model. We begin by introducing the standard clone due to its simplicity [FMT22]. Next, we restate the privacy blanket framework, using consistent terminology with the former [BBGN19]. It aids in identifying the intrinsic connection between the two approaches, which will be explored in Section 4.
@@ 3.1 Standard clone
The intuition behind the standard clone paradigm is as follows [FMT22]: Suppose that $X^{0}$ and $X^{1}$ are neighbouring databases that differ on the first datapoint, $x_{1}^{0}\neq x_{1}^{1}$. A key observation is that for any $\varepsilon_{0}$-DP local randomizer $\mathcal{R}$ and data point $x$, $\mathcal{R}(x)$ can be seen as sampling from the same distribution as $\mathcal{R}(x_{1}^{0})$ with probability at least $e^{-\varepsilon_{0}}/2$ and sampling from the same distribution as $\mathcal{R}(x_{1}^{1})$ with probability at least $e^{-\varepsilon_{0}}/2$. That is, with probability $e^{-\varepsilon_{0}}$ each data point can create a clone of the output of $\mathcal{R}(x_{1}^{0})$ or a clone of $\mathcal{R}(x_{1}^{1})$ with equal probability. Thus $n-1$ data elements effectively produce a random number of clones of both $x_{1}^{0}$ and $x_{1}^{1}$, making it more challenging to distinguish whether the original dataset contains $x_{1}^{0}$ or $x_{1}^{1}$ as its first element.
@@
$\forall\alpha,D_{\alpha}\big{(}f(P)\parallel f(Q)\big{)}\leq D_{\alpha}(P\parallel Q).$
If we can find two probability distributions $P_{0}$ and $P_{1}$ along with a post-processing function $f$ such that $f(P_{0})=\mathcal{M}_{S}(X^{0})$ and $f(P_{1})=\mathcal{M}_{S}(X^{1})$, then it follows that $D_{e^{\varepsilon}}(P_{0}\parallel P_{1})$ is an upper bound for $D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1}))$. We refer to $(P_{0},P_{1})$ as a reduction pair. Different analysis techniques construct different reduction pairs. We first present an intuitive construction of the reduction pair within the standard clone framework, followed by the formal construction.
Definition 5 (Standard Clone Reduction Pair (Intuitive) [FMT22]).
@@
$A_{0}=\begin{cases}0&\text{w.p. }1\\ 1&\text{w.p. }0\\ 2&\text{w.p. }0\end{cases},\ A_{1}=\begin{cases}0&\text{w.p. }0\\ 1&\text{w.p. }1\\ 2&\text{w.p. }0\end{cases},\text{ and }A_{2}=\begin{cases}0&\text{w.p. }\frac{1}{2e^{\varepsilon_{0}}}\\ 1&\text{w.p. }\frac{1}{2e^{\varepsilon_{0}}}\\ 2&\text{w.p. }1-\frac{1}{e^{\varepsilon_{0}}}\end{cases}$
To obtain a sample from $P_{0}$ (or $P_{1}$), sample one copy from $A_{0}$ (or $A_{1}$) and $n-1$ copies of $A_{2}$, the output $(n_{0},n_{1})$ where $n_{0}$ is the total number of 0s and $n_{1}$ is the total number of 1s. Equivalently,
@@
Algorithm 1 Post-processing function of standard clone [FMT22], $f_{1}$
$x_{1}^{0},x_{1}^{1},x_{2},\dots,x_{n}$; $y\in\{0,1,2\}^{n}$
@@
         Let $j$ be a randomly and uniformly chosen element of $[2:n]\setminus J$
@@
An additional observation is that if $\mathcal{R}$ is $\varepsilon_{0}$-DP, then $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ are similar, hence privacy is further amplified [FMT22]. The similarity is characterized by the following lemma:
Lemma 1 ([KOV17]).
@@
With the help of Lemma 1, [FMT22] gives the following decomposition for generic local randomizers:
@@
Theorem 1 (Standard Clone Reduction [FMT22]).
@@
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}^{\mathsf{C}}\parallel P_{1}^{\mathsf{C}}\big{)}$
where $P_{0}^{\mathsf{C}},\ P_{1}^{\mathsf{C}}$ are defined as below (with “$\mathsf{C}$” denoting “standard Clone”) :

### v6

@@ 1 Introduction
Differential Privacy (DP) has become a foundational framework for safeguarding individual privacy while enabling meaningful data analysis [Dwork2006]. In real-world applications, Local Differential Privacy (LDP) is widely adopted as it eliminates the need for a trusted curator by applying noise to each user’s data before aggregation [7, 8, 10, Apple17]. However, this decentralized approach often results in significant utility loss due to excessive noise.
@@
To address this trade-off, the shuffle model introduces a trusted shuffler between users and the aggregator [cheu19, Erlingsson19, Prochlo]. This work considers the single-message shuffle model, where each client sends a report of their data and these reports are then anonymized and randomly shuffled before being sent to the server [feldman2023soda]. Shuffle DP improves the privacy-utility trade-off while minimizing trust assumptions to a single shuffler, making it a promising model for real-world deployment [Wang2020, triangle, cheu22]. For instance, privacy amplification by shuffling was used in Apple and Google’s Exposure Notification Privacy-preserving Analytics [application].
@@
The “amplification-by-shuffling” theorem in the shuffle model implies that when each of the $n$ users randomizes their data using an $\varepsilon_{0}$-LDP mechanism, the collection of shuffled reports satisfies $(\varepsilon(\varepsilon_{0},\delta,n),\delta)$-DP, where $\varepsilon(\varepsilon_{0},\delta,n)\ll\varepsilon_{0}$ for sufficiently large $n$ and not too small $\delta$ [feldman2023soda]. A central theoretical challenge is to characterize and compute this privacy amplification effect. A tighter bound allows for a larger value of $\varepsilon_{0}$ while still achieving $(\varepsilon,\delta)$-DP after shuffling, thereby improving the utility of the mechanism. It also enables a more accurate analysis of the overall $(\varepsilon,\delta)$-guarantees resulting from shuffling $n$ independent $\varepsilon_{0}$-LDP reports.
@@
Researchers are interested in two types of upper bounds on privacy amplification. The first is a generic bound, which provides a uniform guarantee for all $\varepsilon_{0}$-DP local randomizers. The second is a specific bound, which is tailored for a given $\varepsilon_{0}$-DP local randomizer. Generic bounds are primarily studied from a theoretical perspective, as they allow the derivation of asymptotic expressions for $\varepsilon(\varepsilon_{0},n,\delta)$ [Erlingsson19, Balle2019, Feldman2021]. In contrast, specific bounds are usually much tighter and are of significant practical importance, as they provide precise guarantees for concrete mechanisms used in real-world applications.
@@
Among the various strategies proposed to analyze this effect [cheu19, Erlingsson19, checkin], decomposition-based approaches are the most popular [Balle2019, Feldman2021, feldman2023soda]. Two prominent frameworks are the privacy blanket by Balle et al. [Balle2019] and the clone paradigm by Feldman et al., which includes the standard clone [Feldman2021] and the stronger clone [feldman2023soda]. These approaches rely on decomposing the probability distributions induced by a local randomizer under different inputs. A decomposition naturally leads to a reduction, which yields an upper bound on the privacy amplification via shuffling.
@@
The privacy blanket framework designs a tailored decomposition to obtain a bound for each specific local randomizer [Balle2019]. However, this bound cannot be computed precisely in its original form. To address this, the framework resorts to further approximations—such as Hoeffding’s and Bennett’s inequalities—which yield looser but computable bounds. A computable generic bound is also derived in a similar manner.
@@
The standard clone [Feldman2021] adopts a simple, unified decomposition, resulting in a generic bound that applies to all $\varepsilon_{0}$-DP local randomizers. This bound can be computed precisely using a dedicated numerical algorithm and empirically outperforms the computable generic bound derived from the privacy blanket framework. However, the standard clone does not provide specific bounds, except in the special case of $k$-ary Randomized Response.
@@
The stronger clone was introduced with a more refined decomposition and was expected to yield improved bounds for both generic and specific local randomizers [feldman2023soda]. Unfortunately, a critical flaw was later identified in the proof’s core lemma. A corrected version was released on arXiv [feldman2023arxiv], which showed that the original generic bounds hold only for a restricted class of local randomizers. The specific bounds were also revised and replaced with a version that is much weaker and lacks an efficient computation method.
@@ 1.1 Our contributions
In prior work on Shuffle DP, the specific bounds of the blanket framework has seen limited use; instead, its associated analysis has mostly been applied in specific cases when the blanket distribution is uniform, as in $k$-ary randomized response ($k$-$R$) [Balle2019], optimal local hash [Wang2020] or in multi-message shuffle models where uniform blanket noise is directly injected [Luo22]. Although the clone and blanket appear somewhat similar, no formal connection between them has been established.
@@
These two definitions are interconvertible: treating $(D_{1},D_{2},\ldots,D_{k})$ as a single data $D$ yields joint composition as a special case of $k$-fold composition; conversely, setting $D_{1}=D_{2}=\cdots=D_{k}=D$ makes $k$-fold composition a special case of joint composition. Our analysis of joint composition also applies to the $k$-fold composition setting. However, the joint composition framework captures a strictly more general scenario—namely, when $D_{1},D_{2},\dots,D_{k}$ are not necessarily identical and may even belong to different domains. It is important to note that our setting concerns the composition of LDP mechanisms in the single-message shuffle model, in which $D$ or the tuple $(D_{1},D_{2},\dots,D_{k})$ corresponds to the data held by an individual user. Moreover, during the shuffling process, these outputs are treated as a single tuple and are shuffled together as a unit.
Joint composition is widely used in LDP applications such as joint distribution estimation and heavy hitter detection [50, 51, 5, kikuchi2022castellscalablejointprobability]. For example, when each user’s data contains $d$ attributes, applying an $\frac{\varepsilon_{0}}{d}$-LDP mechanism to each attribute ensures overall $\varepsilon_{0}$-LDP while preserving inter-attribute correlations. Our experimental results show that existing methods yield relatively loose bounds in this setting, whereas our algorithm computes significantly tighter results by leveraging the optimal bounds.
@@
It is particularly suited for scenarios in which the analyst aims to compute multiple statistical queries—each corresponding to a distinct local mechanism $\mathcal{R}_{1},\dots,\mathcal{R}_{m}$. Parallel composition partitions the user population into $m$ non-overlapping subsets, assigning each subset to one of the $m$ estimation tasks. As a result, each subset can utilize the full privacy budget $\varepsilon_{0}$, rather than splitting the budget across all tasks. This technique has been widely adopted in the literature [4, 8731512, 10.1145/3299869.3319891, 8758350, 10.1145/3183713.3196906] to achieve better utility compared to naive budget division.
It is worth noting that parallel composition also characterizes Poisson subsampling in the shuffle model [subsample18, steinke2022compositiondifferentialprivacy, Balle_Barthe_Gaboardi_2020]: each user participates in a mechanism $\mathcal{R}^{*}$ with probability $p^{*}$, which is equivalent to the parallel composition $\mathcal{R}_{[1,2]}^{\parallel,\boldsymbol{p}}(x)$, where $\mathcal{R}_{1}(x)=\mathcal{R}^{*}(x)$, $\mathcal{R}_{2}(x)=\bot$, and $p_{1}=p^{*}$, $p_{2}=1-p^{*}$.
To compute the privacy amplification of parallel composition in the shuffle model was previously studied in [tv_shuffle], but their analysis inherited a technical flaw from [feldman2023soda], resulting in incorrect conclusions. Our method provides the first correct computation of the optimal bound in this setting.
@@ 1.2 Related Work
We now discuss prior work most relevant to the technical contributions of our paper. Koskela et al. [koskela2023numerical] investigate $k$-fold composition of Shuffle DP protocols. Their approach computes the Privacy Loss Random Variable (PLRV) of a dominating pair for a single-round shuffle protocol, and then applies the FFT to analyze multiple rounds (i.e., composition over $k$ executions) [koskela20b]. In contrast, our work focuses on joint composition, where a single-round shuffle is applied to a composite local randomizer constructed from multiple sub-randomizers.
Some subtle aspects warrant further clarification, particularly regarding how coupling among sub-randomizers affects the composition analysis. In the single-message shuffle model, when the local randomizer is a joint composition, the outputs of the sub-local randomizers for a single user are inherently coupled. From a theoretical standpoint, privacy analysis in this setting ($\mathcal{S}\circ(\mathcal{M}_{1}\times\mathcal{M}_{2}\times\dots\times M_{k})^{n}$) is not equivalent to applying composition to the individual mechanisms $\mathcal{R}_{i}$ after shuffling $(\mathcal{S}_{1}\circ\mathcal{M}_{1}^{n},\mathcal{S}_{2}\circ\mathcal{M}_{2}^{n},\dots,\mathcal{S}_{k}\circ\mathcal{M}_{k}^{n})$. Furthermore, the standard compositional properties of the PLRV do not apply.
In [koskela2023numerical], computing the PLRV requires a dominating pair [Zhu22], obtained either from the stronger clone (in the general case) or from the privacy blanket (in the case of $k$-$R$). In the general case, [koskela2023numerical] inherits a flaw that originates from the original version of the stronger clone construction (see Section 3.3). For any specific local randomizer, our work demonstrates that the privacy bound derived from the blanket framework is strictly tighter than that obtained via the clone.
Theoretically, computing the PLRV of $(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$ (as defined in Section 3.2) and applying the Hockey-Stick divergence $D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$ yields the same amplification bound as computing it directly from the GPARV; both approaches evaluate $D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$. However, for general local randomizers, the PLRV of $(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$ is analytically intractable due to the complex structure of the blanket distribution. GPARV overcomes this limitation by expressing the privacy loss in a convolution-friendly form, enabling efficient FFT-based evaluation even in the absence of a closed-form expression for the PLRV.
Subsequent work.
A recent paper to appear at CSF 2026 extends our decomposition framework to the analysis of re-identification attacks in the shuffle model, showing that the same hierarchy of decompositions—where the blanket decomposition remains optimal—also holds in this context [Su26CSF].
@@ 1.3 Roadmap
Section 5 introduces the Generalized Privacy Amplification Random Variable (GPARV) and an FFT-based algorithm for computing the optimal privacy amplification bound. We also clarify the relationship between GPARV and the original PARV introduced by Balle et al. [Balle2019]. Section 6 presents our experimental results.
@@ 2 Preliminaries
Following conventions in the shuffle model based on randomize-then-shuffle [Balle2019, cheu19], we define a single-message protocol $\mathcal{P}$ in the shuffle model as a pair of algorithms $\mathcal{P}=(\mathcal{R},\mathcal{A})$, where $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$, and $\mathcal{A}:\mathbb{Y}^{n}\to\mathbb{O}$. We call $\mathcal{R}$ the local randomizer, $\mathbb{Y}$ the message space of the protocol, $\mathcal{A}$ the analyzer, and $\mathbb{O}$ the output space.
@@ 3 Review of Existing Analysis Techniques
In this section, we review existing analysis techniques for studying privacy amplification in the shuffle model. We begin by introducing the standard clone due to its simplicity [Feldman2021]. Next, we restate the privacy blanket framework, using consistent terminology with the former [Balle2019]. It aids in identifying the intrinsic connection between the two approaches, which will be explored in Section 4.
@@ 3.1 Standard clone
The intuition behind the standard clone paradigm is as follows [Feldman2021]: Suppose that $X^{0}$ and $X^{1}$ are neighbouring databases that differ on the first datapoint, $x_{1}^{0}\neq x_{1}^{1}$. A key observation is that for any $\varepsilon_{0}$-DP local randomizer $\mathcal{R}$ and data point $x$, $\mathcal{R}(x)$ can be seen as sampling from the same distribution as $\mathcal{R}(x_{1}^{0})$ with probability at least $e^{-\varepsilon_{0}}/2$ and sampling from the same distribution as $\mathcal{R}(x_{1}^{1})$ with probability at least $e^{-\varepsilon_{0}}/2$. That is, with probability $e^{-\varepsilon_{0}}$ each data point can create a clone of the output of $\mathcal{R}(x_{1}^{0})$ or a clone of $\mathcal{R}(x_{1}^{1})$ with equal probability. Thus n1n-1 data elements effectively produce a random number of clones of both $x_{1}^{0}$ and $x_{1}^{1}$, making it more challenging to distinguish whether the original dataset contains $x_{1}^{0}$ or $x_{1}^{1}$ as its first element.
@@
$D_{\alpha}\big(f(P)||f(Q)\big)\leq D_{\alpha}(P||Q).$
If we can find two probability distributions $P_{0}$ and $P_{1}$ along with a post-processing function $f$ such that $f(P_{0})=\mathcal{M}_{S}(X^{0})$ and $f(P_{1})=\mathcal{M}_{S}(X^{1})$, then it follows that $D_{e^{\varepsilon}}(P_{0},P_{1})$ is an upper bound for $D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1}))$. We refer to $(P_{0},P_{1})$ as a reduction pair. Different analysis techniques construct different reduction pairs. We first present an intuitive construction of the reduction pair within the standard clone framework, followed by the formal construction.
Definition 5 (Standard Clone Reduction Pair (Intuitive) [Feldman2021]).
@@
[…]
To obtain a sample from $P_{0}$ (or $P_{1}$), sample one copy from $A_{0}$ (or $A_{1}$) and n1n-1 copies of $A_{2}$, the output $(n_{0},n_{1})$ where $n_{0}$ is the total number of 0s and $n_{1}$ is the total number of 1s. Equivalently,
@@
Algorithm 1 Post-processing function of standard clone [Feldman2021], $f_{1}$
  Meta parameter: $x_{1}^{0},x_{1}^{1},x_{2},\dots,x_{n}$
$y\in\{0,1,2\}^{n}$
@@
   Let $j$ be a randomly and uniformly chosen element of $2:n]\setminus J$
@@
An additional observation is that if $\mathcal{R}$ is $\varepsilon_{0}$-DP, then $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ are similar, hence privacy is further amplified [Feldman2021]. The similarity is characterized by the following lemma:
Lemma 1 ([KOV15]).
@@
With the help of Lemma 1, [Feldman2021] gives the following decomposition for generic local randomizers:
@@
Theorem 1 (Standard Clone Reduction [Feldman2021]).
@@
$D_{e^{\varepsilon}}\big(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big)\leq D_{e^{\varepsilon}}\big(P_{0}^{\mathsf{C}},P_{1}^{\mathsf{C}}\big)$
where $P_{0}^{\mathsf{C}},\ P_{1}^{\mathsf{C}}$ are defined as below (with $\mathsf{C}$” denoting “standard Clone”) :